OpenStudy (anonymous):

Madeline was completing the square and her work is shown below. Identify the line where she made her mistake. f(x) = 3x^2 + 6x − 7 Line 1: f(x) = 3(x + 2x) − 7 Line 2: f(x) = 3(x + 2x + 1) − 7 − 1 Line 3: f(x) = 3(x + 1)2 − 8

OpenStudy (anonymous):

Where did the ^2 go?

OpenStudy (anonymous):

f(x) = 3x^2 + 6x − 7 Line 1: f(x) = 3(x + 2x) − 7 Line 2: f(x) = 3(x + 2x + 1) − 7 − 1 Line 3: f(x) = 3(x + 1)^2 − 8

OpenStudy (anonymous):

the ^2 on the second line is missing should be 3(x^2+2x)-7

OpenStudy (anonymous):

well in line two should she have subtracted one on the end? I assume the ^2 is in every line

OpenStudy (anonymous):

she sould have subtracted 3

OpenStudy (anonymous):

yep!!

OpenStudy (anonymous):

assuming he messed up ^2 the answer is line 2